OpenStudy (anonymous):

A cube has a surface area of 294 square cm. Find the length of one side

OpenStudy (anonymous):

the surface area of a cube =6*(length of each side)^2

OpenStudy (akashdeepdeb):

Hey, @jera25617 To solve this question you need to understand the cube! It looks like this:  Alright, so that was the basics. Now, you are given that the Total surface area = \(6s^2 = 294\) Thus, after solving that for 's', we get \(s^2=49 ~~~~~=> ~~~~s = 7\) Thus, your edge length og the cube = 7cm! :D Understood? :)